An automobile battery manufacturer claims that its midgrade battery has a mean life of 50 months with a standard deviation of 6 months. Suppose the distribution of battery lives of this particular brand is approximately normal, find the probability that a randomly selected battery of this type will last less than 48 months.

1
Expert's answer
2022-03-23T16:54:21-0400

"P(X<48)=P(Z<\frac{48-50}{6})=P(Z<-0.33)=0.3707."